OpenStudy (lifeisadangerousgame):

2/3 * ln(x-3)^6 + 1/4[ln x^-2 - ln(x^2 - 2x - 3)^2] Can someone solve this? I have done the problem already but I think the way I'm doing it now, is different than how I did it two days ago.

OpenStudy (lifeisadangerousgame):

\[\frac{ 2 }{ 3 }\ln (x-3)^{6} + \frac{ 1 }{ 4}\left[ \ln (x ^{-2}) - \ln (x^2 - 2x -3)^2 \right]\]

OpenStudy (mathmale):

Where are the instructions? You can't really "solve" this expression because it's not an equation. Only equations can be "solved."